The following give the number of bits in an encryption key. For each, calculate the number of trials, on average, it will take to crack that key in powers of ten. a. 40 bits b. 56 bits c. 64 bits d. 80 bits e. 128 bits f. 256 bits
Answer #1

The number of trial is found by the following algebraic equation:

log( 2number of bits )= 10x trails

a. 40 bits

= log(240) = 12.0411998266

In power of 10 it is rounded to 1012.

40 = rounded 1012 Trials.

b. 56 bits

= log(256) = 16.8576797572

In power of 10 it is rounded to 1016.

56 = rounded 1016 Trials.

c. 64 bits

= log(264) = 19.2659197225

In power of 10 it is rounded to 1019.

64 = rounded 1019 Trials.

d. 80 bits

= log(280) = 24.0823996531

In power of 10 it is rounded to 1024.

80 = rounded 1024 Trials.

e. 128 bits

= log(2128) = 38.531839445

In power of 10 it is rounded to 1038.

128 = rounded 1038 Trials.

e. 256 bits

= log(2256) = 77.06367889

In power of 10 it is rounded to 1077.

256 = rounded 1077 Trials.